### v1

@@ 6 Conclusion
Specifically, we focus on two algorithms, $\mathit{odt}_{\text{vec}}$ and $\mathit{odt}_{\text{rec}}$—the former offering superior hardware compatibility, while the latter achieving stronger theoretical efficiency. We provide comprehensive experimental results analyzing the computational efficiency of $\mathit{odt}_{\text{vec}}$ and $\mathit{odt}_{\text{rec}}$. Additionally, we conduct a thorough assessment of the generalization capabilities of hypersurface decision tree models and axis-parallel decision tree algorithms, including those learned by CART and Brița et al. (2025)’s ConTree algorithm, using both synthetic and real-world datasets.

### v2

@@ 6 Conclusion
Specifically, we focus on two algorithms, $\mathit{odt}_{\text{vec}}$ and $\mathit{odt}_{\text{rec}}$—the former offering superior hardware compatibility, while the latter achieving stronger theoretical efficiency. We provide comprehensive experimental results analyzing the computational efficiency of $\mathit{odt}_{\text{vec}}$ and $\mathit{odt}_{\text{rec}}$. Additionally, we conduct a thorough assessment of the generalization capabilities of hypersurface decision tree models and axis-parallel decision tree algorithms, including those learned by CART and Brita et al. (2025)’s ConTree algorithm, using both synthetic and real-world datasets.